Step 1. Open Excel and complete first two columns with any numbers

Make 7-8 rows.

Type below first column "R = "

We will use Excel function CORREL(Array-1, Array-2).

STEP 2.

Click on cell B11 and type: =CORREL(

Highlight all numbers in A-column starting with cell A2 (do not include title).

That will be Array-1 (input range).

Type comma after Array-1 and highlight Array-2 from B-column.

Press ENTER. You will see in cell B11 Coefficient of Linear Correlation.

STEP 3. Analyze the R-value. Does it mean we have strong positive,

strong negative, weak correlation or none linear correlation.







Part 2. Apply t-Distribution to construct Confidence Interval for Population Mean (Chapter 10).

STEP 1. Open Excel and write any 7 numbers that represents Sample results.

STEP 2. Use Excel functions Average() and STDEV() to find sample mean <x>

and sample standard deviation (s).

STEP 3. Use the Table for t-Distribution Table t-Distribution.pdf to find t-value for row df= n-1 = 7-1 = 6

and column 0.05 in two tails (this will be 95% Confidence Interval).

Calculate the Margin of Error: E = t*s/√(n)

STEP 4. Write Confidence Interval for Population Mean in format:

<x> - E, <x> + E
Answer #1

A C D E F G H I J K L M N Y 215 Correlation between X and y is, CORREL(X,Y) 185 r= 0.987238 406 Here from ther value we conclIVIN 13 PART II = STEP STEP II AVERAGE(X) 16.42857 STDEV(X) 3.598942 STEP III Critical Value t(n-1), a/2 = 1/7-1,0.05/2) = 16